A 65.0-kg runner has a speed of 5.20 m/s at one instant during a long-distance

event. (a) What is the runner’s kinetic energy at this instant? (b) How much net

work is required to double his speed?


1
Expert's answer
2022-05-30T15:17:27-0400

(a)

"E_{k1}=\frac{mv_1^2}{2}=\frac{65.0*5.20^2}{2}=879\:\rm J"

(b)

"W=E_{k2}-E_{k1}=4E_{k1}-E_{k1}=3E_{k1}\\
=3*879=2640\:\rm J"